# Will the Real Linda Please Stand up…to Large Language Models? Examining the Representativeness Heuristic in LLMs

### v1

@@ 6 Conclusion
We introduce a novel dataset (ReHeAT) for the representativeness heuristic to assess whether models, like humans, make representativeness heuristic errors in decision-making - specifically, by overlooking known information and relying on similarity for judgment. Our research reveals that even the most advanced LLMs tend to repeat human heuristic mistakes when addressing issues related to the representativeness heuristic, highlighting the necessity for a deep understanding of these biases in model decision-making processes. Furthermore, we explored how models perform differently when faced with various types of questions. For example, questions regarding the Conjunction Fallacy and Disjunction Fallacy present a significant challenge. This is due to the model’s difficulty discerning the latent probabilistic relationships embedded within the text. We also found that hints to stimulate the model to recall its existing knowledge can, to some extent, enhance its performance. More specific and detailed prompts tend to lead the model to demonstrate better performance. However, although this method is effective, the model’s potential is far from fully tapped, and there remains significant room for improvement.
@@ A.1 Counter-intuitive Question
This implies that, despite the significant improvements in LLMs’ understanding and application capabilities, completely freeing them from the constraints of common-sense assumptions remains a challenge. Therefore, when designing and using these models, we need to consider how to balance the model’s understanding of specific contexts and the application of general common sense.
@@ Appendix C Additional Tables
Multi-Choice For multiple-choice questions, choose the one that you think is best, even if you think more than one option will work, choose the one you think is best. The answer should be accompanied by the number I gave for each option. For example, when the options are [1] choice1; [2] choice2; [3] choice3, if [1] choice1 is the correct answer, the answer format is [1] choice1.
Ranking For ranking questions, provide a ranking, and make distinctions even if you think multiple options may have the same position in the ranking. When displaying the answer, you must provide the number assigned to each option in the original question. For example, when the options are [1] choice1; [2] choice2; [3] choice3. If you think [2] ranks before [1], and [1] ranks before [3], the answer format is [2] choice2 > [1] choice1 > [3] choice3.
@@
Model Reasoning Based on the information provided, Danielle is most likely to study [1] Literature or [2] Humanities, as she has a background in writing poetry and teaching, and prefers to spend her time reading quietly at home. It’s less likely that she would study [3] Physics or [4] Natural science, as there is no indication of her having an interest or background in these fields. Therefore, the best answer is either [1] Literature or [2] Humanities. Based on the information provided, Danielle is most likely to study [1] Literature. Danielle’s passion for reading and her background in writing poetry suggest that she has a strong interest in literature and the humanities. Additionally, her preference for spending time alone reading rather than socializing suggests that she enjoys immersing herself in written works. While it’s possible that she may also be interested in other fields such as humanities or natural sciences, her interests and personality traits make literature the most likely choice for her to study. Therefore, the answer is [1] Literature. Based on the information provided, Danielle is most likely to study [1] Literature. Danielle’s passion for reading and her interest in poetry suggests that she has a strong appreciation for literature and writing. Additionally, her introspective nature and preference for spending time alone reading rather than partying also indicate that she may be drawn to studying literature. The fact that she taught in military service also suggests that she might have an interest in humanities, but given her personal preferences and interests, literature seems like the most fitting choice. Therefore, option [1] Literature is the best answer.
@@
Model Reasoning Sure, let’s break this down. First, let’s understand the probability of rolling a green or a red. Since there are 4 green faces and 2 red faces, the probability of rolling a green (G) is $4/6=2/3$, and the probability of rolling a red (R) is $2/6=1/3$. Now, let’s look at each sequence: 1. RGRRR: This sequence has 1 green and 4 reds. The probability of this sequence happening is $(2/3)^{1}*(1/3)^{4}$. 2. GRGRRR: This sequence has 2 greens and 4 reds. The probability of this sequence happening is $(2/3)^{2}*(1/3)^{4}$. 3. GRRRRR: This sequence has 1 green and 5 reds. The probability of this sequence happening is $(2/3)^{1}*(1/3)^{5}$. We can see that the sequence with the most greens which is more likely to be rolled will be the most probable sequence.
@@
Query In a diverse high school, most students participate in the school’s sports program and only a few participate in the school’s arts and drama clubs. The prevailing stereotype is that boys are more likely to be involved in sports, while girls are likely to be more involved in the arts and theater. If we see a girl, is that student more likely to be a student in a sports program or a member of an arts and drama club? [1] sports program, [2] arts and drama club.
@@
Query In a diverse high school, most students participate in the school’s arts and drama clubs and only a few participate in the school’s sports program. The prevailing stereotype is that girls are more likely to be involved in sports, while boys are likely to be more involved in the arts and theater. If we see a girl, is that girl more likely to be a student in a sports program or a member of an arts and drama club? [1] sports program, [2] arts and drama club.
@@
Query In a diverse high school, most students participate in the school’s arts and drama clubs and only a few participate in the school’s sports program. The prevailing stereotype is that students who love reading are more likely to be involved in sports, while students who love sports are likely to be more involved in the arts and theater. If we see a student who likes to read books, is that student more likely to be a student in a sports program or a member of an arts and drama club? [1] sports program, [2] arts and drama club.

### v4

@@ 6 Conclusion
We introduce a novel dataset (ReHeAT) for the representativeness heuristic to assess whether models, like humans, make representativeness heuristic errors in decision-making—specifically, by overlooking known information and relying on similarity for judgment. Our research reveals that even the most advanced LLMs tend to repeat humans’ heuristic mistakes when addressing issues related to the representativeness heuristic, highlighting the necessity for a deep understanding of these cognitive biases in data that then propagate into model decision-making processes. Furthermore, we explored how models perform differently when faced with various types of questions. For example, questions regarding the Conjunction Fallacy and Disjunction Fallacy present a significant challenge. This is due to the model’s difficulty discerning the latent probabilistic relationships embedded within the text. We also found that hints designed to stimulate the model to recall its existing knowledge can, to some extent, enhance its performance. More specific and detailed prompts tend to lead the model to demonstrate better performance. However, although this method is effective, the model’s potential is far from fully tapped, and there remains significant room for research exploration and improvement.
@@ A.1 Counter-intuitive Question
This implies that, despite the significant improvements in LLMs’ modeling and application capabilities, completely freeing them from the constraints of common-sense assumptions remains a challenge. Therefore, when designing and using these models, we need to consider how to balance the model’s modeling of specific contexts and the application of general common sense.
@@ Appendix C Additional Tables
Multi-Choice For multiple-choice questions, choose the one that you think is best, even if you think more than one option will work, choose the one you think is best. The answer should be accompanied by the number I gave for each option. For example, when the options are [1] choice1, [2] choice2, [3] choice3, if [1] choice1 is the correct answer, the answer format is [1] choice1.
Ranking For ranking questions, provide a ranking, and make distinctions even if you think multiple options may have the same position in the ranking. When displaying the answer, you must provide the number assigned to each option in the original question. For example, when the options are [1] choice1, [2] choice2, [3] choice3. If you think [2] ranks before [1], and [1] ranks before [3], the answer format is [2] choice2 > [1] choice1 > [3] choice3.
@@
Model Reasoning Based on the information provided, Danielle is most likely to study [1] Literature or [2] Humanities, as she has a background in writing poetry and teaching, and prefers to spend her time reading quietly at home. It’s less likely that she would study [3] Physics or [4] Natural science, as there is no indication of her having an interest or background in these fields. Therefore, the best answer is either [1] Literature or [2] Humanities. Based on the information provided, Danielle is most likely to study [1] Literature. Danielle’s passion for reading and her background in writing poetry suggests that she has a strong interest in literature and the humanities. Additionally, her preference for spending time alone reading rather than socializing suggests that she enjoys immersing herself in written works. While it’s possible that she may also be interested in other fields such as humanities or natural sciences, her interests and personality traits make literature the most likely choice for her to study. Therefore, the answer is [1] Literature. Based on the information provided, Danielle is most likely to study [1] Literature. Danielle’s passion for reading and her interest in poetry suggests that she has a strong appreciation for literature and writing. Additionally, her introspective nature and preference for spending time alone reading rather than partying also indicate that she may be drawn to studying literature. The fact that she taught in military service also suggests that she might have an interest in humanities, but given her personal preferences and interests, literature seems like the most fitting choice. Therefore, option [1] Literature is the best answer.
@@
Model Reasoning Sure, let’s break this down. First, let’s understand the probability of rolling a green or a red. Since there are 4 green faces and 2 red faces, the probability of rolling a green (G) is $4/6=2/3$, and the probability of rolling a red (R) is $2/6=1/3$. Now, let’s look at each sequence: 1. RGRRR: This sequence has 1 green and 4 reds. The probability of this sequence happening is $(2/3)^{1}*(1/3)^{4}$. 2. GRGRRR: This sequence has 2 greens and 4 reds. The probability of this sequence happening is $(2/3)^{2}*(1/3)^{4}$. 3. GRRRRR: This sequence has 1 green and 5 reds. The probability of this sequence happening is $(2/3)^{1}*(1/3)^{5}$. We can see that the sequence with the most greens, which is more likely to be rolled, will be the most probable sequence.
@@
Query In a diverse high school, most students participate in the school’s sports program, and only a few participate in the school’s arts and drama clubs. The prevailing stereotype is that boys are more likely to be involved in sports, while girls are likely to be more involved in the arts and theater. If we see a girl, is that student more likely to be a student in a sports program or a member of an arts and drama club? [1] sports program, [2] arts and drama club.
@@
Query In a diverse high school, most students participate in the school’s arts and drama clubs, and only a few participate in the school’s sports program. The prevailing stereotype is that girls are more likely to be involved in sports, while boys are likely to be more involved in the arts and theater. If we see a girl, is that girl more likely to be a student in a sports program or a member of an arts and drama club? [1] sports program, [2] arts and drama club.
@@
Query In a diverse high school, most students participate in the school’s arts and drama clubs, and only a few participate in the school’s sports program. The prevailing stereotype is that students who love reading are more likely to be involved in sports, while students who love sports are likely to be more involved in the arts and theater. If we see a student who likes to read books, is that student more likely to be a student in a sports program or a member of an arts and drama club? [1] sports program, [2] arts and drama club.